OpenStudy (anonymous):

using the quadratic formula 3x^2-4x=2

OpenStudy (zepp):

Set one side to 0, so it would be 3x^2-4x-2 = 0 And then you can use the quadratic formula. :)

hero (hero):

Hint: The quadratic formula is: \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] And in this case the values of the coefficients are a = 3 b = -4 c = -2

OpenStudy (anonymous):

so it would be 4 pulse or minus sqrt 40 over 2?

hero (hero):

So input these values into the formula: \[x = \frac{4 \pm \sqrt{(-4)^2 - 4(3)(-2)}}{2(3)}\] then simplify

OpenStudy (zepp):

It isn't 2, the denominator is 2a, you forgot something there. It should be 6

hero (hero):

zepp, you should probably make it clear who you are directing your comment towards. I simply wrote the formula, then inputted the values of a,b, and c.

hero (hero):

I wrote "then simplify" because it needs to be simplified.

OpenStudy (zepp):

I'm sorry, would be more careful next time. :)